OpenStudy (jchick):

Which of the following is not an equivalent form of the compound inequality x + 15 less than or greater to 12 and x + 15 greater than or equal to 5

OpenStudy (jchick):

a. 12 greater than or equal to x + 15 greater than or equal to 5 b. -3 greater than or equal to x greater than or equal to -10 c. A number line with a closed circle on -10, a closed circle on -3, and shading in between. d. A number line with a closed circle on -10, shading to the left, and a closed circle on -3, shading to the right.

OpenStudy (jchick):

I was thinking b

OpenStudy (jchick):

@Nnesha

OpenStudy (jhannybean):

For the first part did you mean less than or equal to? I dont understand the `less than or greater to`

OpenStudy (jchick):

Yes I did

OpenStudy (jchick):

@Nnesha

OpenStudy (jhannybean):

Well ok, then \[x+15 \le 12 \implies x \le ~?\]\[5 \le x+15 \implies x \le ~?\]

OpenStudy (jchick):

x<=-3

OpenStudy (jhannybean):

solve for both.

OpenStudy (jchick):

x>=-10

OpenStudy (jchick):

I was just working on the second one

OpenStudy (jhannybean):

Alright, so you have \[1.~ x\le -3 ~,~ 2. -10\le x\] We can combine 1. and 2. to form one compound inequality as so: \[-10 \le x \le-3\]

OpenStudy (jhannybean):

One sec, I will brb.

OpenStudy (jchick):

Ok thanks

OpenStudy (jchick):

U still there?

OpenStudy (jchick):

Answer is D right?

OpenStudy (jchick):

I am lost now.........

OpenStudy (jhannybean):

Yes it is.

OpenStudy (jhannybean):

Can you explain to me why you think it is D?

OpenStudy (jchick):

I think it was d because it was the only one that didn't match up when I graphed it However I wasn't sure that I was graphing it correctly